# references.bib

### v1

@@ 4.1.1 AI Assistants Interacting with E2EE Data
As discussed in Section 2.2, a model’s training data consists of vast amounts of information which can be continuously supplemented by data collected from users (e..g, queries, usage patterns) after the model is deployed and used for inference. This data would necessarily be plaintext as a model cannot be trained on encrypted text. As such, if user data is collected and incorporated into the training process, the model is inherently, itself, a derivative of E2EE content.
Additionally, models have been observed to reproduce portions of the training data in their outputs, at times producing significant portions of training data, sometimes verbatim or nearly verbatim [cooper2024files, petroni2019languagemodelsknowledgebases]. The exact reasons behind this behavior are not fully understood, but it is widespread and believed by many experts to be a fundamental limitation of current machine learning techniques, and unlikely to be eliminated in the near term or without a significant shift in the underlying approaches to AI development. As a result, violating the E2EE guarantee by training AI assistants on plaintext user data poses a tangible risk to the security of user information, as models may expose information from the training data during regular usage of the model, thereby leaking sensitive information.
The memorization and reproduction of training data in models also introduces the risk of individuals exploiting this behavior to extract certain data points and patterns. Prior work has shown that this can be done to expose information in the training data, which would additionally compromise the confidentiality of E2EE data [274574]. In particular, individuals may perform adversarial attacks, involving using the model (often in a black-box manner) as a means to acquire private data. Adversarial attacks are an active research area, and come in various forms, including membership inference attacks, where an adversary can determine whether a particular data point is part of the model’s training set [shokri2017membership, hu2022membership]; and training data extraction attacks, where an adversary can recover elements of the model’s training data often by generating candidate data points and using membership inference to verify whether the items are in the training set [274574].
@@ 4.1.2 Key Considerations
The risks outlined above underscore the following critical considerations.
@@
    Which parties have access to the inputs and outputs from the model? Intermediaries that process plaintext user queries before relaying them to the model provider may have access to inputs and outputs from the model, in the form of chat logs. In order to preserve the guarantees of E2EE, inputs and outputs that depend on user messages should only be visible to the intended recipients of the messages.
@@ 4.2 How Privacy Technologies Can and Cannot Help
The privacy concerns outlined in the preceding subsection have long been acknowledged by the machine learning, privacy, and security communities. As a result, there exist various techniques to enhance privacy—according to various metrics—during inference and training. In this section, we evaluate whether the types privacy protection that these techniques offer meet E2EE confidentiality.
We introduce a comprehensive framework which identifies the key design considerations for implementing AI assistants in E2EE environments. Our framework can be used to evaluate and compare different implementations of AI assistants (which incorporate various privacy techniques), and the degree to which they uphold E2EE security. These considerations are interdependent: at each consideration point, specific design choices can either affirm the configuration’s compliance with E2EE or breach it (thereby rendering further considerations irrelevant). Figure 5 summarizes this evaluation framework.
@@
Based on our analysis, we evaluate the ways in which these configurations do and do not mitigate the security concerns associated with AI training and inference, and whether they are compatible with E2EE. While some of the configurations reflect existing practices, others represent novel configurations. In our analysis, we include the following privacy-enhancing techniques: endpoint-local training and inference, fully homomorphic inference (FHE), trusted execution environments (TEEs), data sanitization, and differential privacy. Our framework can be generalized beyond these examples.
@@
The first design parameter is the location where data is processed, i.e., where the model is hosted. We distinguish between endpoint-local and non-endpoint-local designs, using the definitions from Section 3.
@@
Using endpoint-local models is not simple, however, as most consumer grade devices are unable to support the hefty computations involved in running an AI assistant. In some circumstances, local processing is feasible: e.g., in Apple Intelligence, as discussed in Section 2.5. But even with the benefits of Apple’s custom hardware, not all AI assistant queries can be supported with the limited computational faculties of a smartphone, and certain features require sending the message in plaintext form off the device.
When non-endpoint-local models are used to answer queries, the compatibility with E2EE depends on the further design considerations #2 and #3 given subsequently.
Consideration #2: If the model is non-endpoint local, is the model “protected” or not?
For non-endpoint local models, the privacy of data depends on whether the model is stored in plaintext or protected (i.e., encrypted or contained in secure hardware) form in the servers. Plaintext inference and training is not compatible with E2EE, as they would expose user data to third parties. Thus the only viable option is if data is “protected” in some way that makes it inaccessible to third parties, while still allowing certain computations (inference and training) over this data. There are various technologies that could implement this functionality—such as fully homomorphic encryption (FHE), and trusted execution environments (TEEs)—which differ on their confidentiality guarantees, performance, and other practical considerations. The compatibility of protected models with E2EE depends on the nature of the confidentiality offered by the technology being used, which is the focus of consideration #3.
Consideration #3: If the model is non-endpoint local and is protected, what are the type of the confidentiality guarantees?
As discussed in Section 2.4.2, different security technologies—such as TEEs and encryption—provide different types of confidentiality under different conditions.
Established cryptographic techniques tend to provide the same type of security as E2EE under types of conditions, and are therefore more likely to be compatible with E2EE. Fully homomorphic encryption (FHE) provides a good example. FHE allows a party to perform arbitrary computations on encrypted data without first needing to decrypt it, with the result of the computation also obtained in encrypted form. FHE makes it possible to “outsource” a computation on a piece of private data without revealing anything about the data to the outsourcing provider. FHE would thus allow non-endpoint-local models to process user messages for inference without decrypting them—at least in theory.
@@
As described in Section 2.4, hardware-backed security is different from E2EE security. Hardware solutions represent a substantial security/privacy improvement over plaintext inference, which may well be a fruitful avenue for exploration, beyond the scope of this paper. Our scope is to stress that the importance of clearly distinguishing them from E2EE and cryptographic security guarantees.
@@
Regardless of all of the above, training introduces an additional critical consideration: which users can query the models trained with the E2EE data of which users? Since the model is a derivative of its training data, it is important that queries over this training data are consistent with the guarantees of E2EE.
There are two possible configurations for this design choice: shared or per-user models, as defined in Section 3. In the former, E2EE data is used to train a model that users other than the sender or the recipient can use for inference (for example, the application could maintain a single server-side model, which is trained on the E2EE data of all users and that all user devices can query). Conversely, in the latter, users can only query a model that is trained exclusively on their own E2EE data, and this data is not used to train any other models.
@@
Shared models are not compatible with E2EE, since users would be able to receive inference outputs that are a function of other users’ E2EE data, which violates E2EE confidentiality. In practice, this means that E2EE data may be exposed to entities that can query the model (e.g., due to the possibility of leakage of training data and adversarial attacks). As we discuss in more detail in Section 4.2.1, mitigations for adversarial attacks (e.g., differential privacy) are insufficient to fully address this concern. Shared models are incompatible with training on E2EE content, irrespective of the implementation.
@@
On the other hand, per-user models can be compatible with E2EE, provided that applications enforce strict isolation between user data, and that any components that depend on E2EE data must themselves be “protected” in the sense outlined in considerations #2 and #3. There are a number of ways an application could support per-user models. For example, a service could maintain a separate copy of a foundation model for each user in the system, which is fine-tuned with that user’s data. This approach would maintain strict isolation, but require large amounts of resources for applications with many users. A more practical approach is to use retrieval augmentation techniques: the application hosts a single foundation model, which is not trained on user data, but is dynamically refined during inference using user-dependent data sources [martineau_what_2023]. [^33] See [wutschitz2023rethinking] for more on retrieval augmentation. That is, the application would store E2EE data for every user (e.g., locally) [martineau_what_2023], and then, during inference, the data for the querying user is retrieved and incorporated into the model in some way, which temporarily fine-tunes the model. After the query is processed, the base model is reverted back to its original state, which was not dependent on any E2EE data, resulting in a stateless augmentation process. While this reversion or statelessness would largely address the specific concern of Consideration #4—namely, which users can query the models trained with the E2EE data of which users—the storage and processing of the per-user E2EE data and the temporarily fine-tuned model (which is itself E2EE data) would, in turn, need to be scrutinized under Considerations #1–#3 and comply with the recommendations we put forward throughout this paper.
@@
l—llll Implementation Location Protected Guarantees Shared or per-user
@@
Table 1: Examples of how different implementations of AI assistants map to our design framework. Green tags denote an implementation compatible with E2EE, yellow tags denote an implementation that is not compatible with E2EE but has additional protections, and red tags denote implementations that are most vulnerable.
4.2.1 A Note on Privacy-Preserving Training
A number of techniques have been put forth towards mitigating adversarial attacks, such as differential privacy (DP) and data sanitization.
@@ 4.2.1 A Note on Privacy-Preserving Training
By definition, statistics computed on datasets reveal some information about the contents of the data. Differential privacy is a method for designing algorithms to compute statistics over datasets of personal information, in such a way that the results are guaranteed not to depend “too much” on any individual data entry. In contrast with encryption, which offers a binary type of security—a ciphertext must hide all information about the underlying plaintext—a key feature of DP is that it is parameterized by an “acceptable” amount of so-called “privacy loss” which represents dependence upon individual data entries (encoded through a parameter $\epsilon$[DBLP:conf/tamc/Dwork08]. DP’s configurable privacy loss parameter stands in sharp opposition to the strict confidentiality offered by encryption. Applying DP to machine learning specifically is an area of active research. Training with DP incurs computational overhead and loss in accuracy compared to a non-DP model—the less the acceptable privacy leakage, the greater the overhead and the loss [nist_dp_ml, Ponomareva_2023].
@@
Data sanitization can mitigate attacks of the kind mentioned in Section 4.1.1 by removing particularly “sensitive” information from messages before adding them to a training dataset for an AI model. This of course presents the difficulty of determining what information within a message is sensitive; sensitivity depends on context and there is no generally correct heuristic for performing this procedure [brown2022does]. Even putting aside this difficulty, the focus on “sensitive” components of a message stands in sharp opposition to the strict confidentiality offered by encryption which is agnostic to the nature of message contents—and indeed provides additional security benefits by virtue of protecting mundane communications alongside sensitive ones (as discussed in Section 2.1.2).
While privacy-preserving training techniques such as the above have made important steps towards mitigating the feasibility of attacks, they do not meet the security guarantees of E2EE: they reduce information leakage in training data, but they do not prevent it completely—and their preventative measures are not based on the security of encryption, so the protection they provide is of a different nature to E2EE security.
The guarantees of private training techniques are contextual, depending on factors such as the type and structure of data, the number of data points, etc. Conversely in E2EE, confidentiality holds for any possible message under any context. Thus, the incompatibility of private training and E2EE is not a matter of computational limitations, but rather a definitional one. In other words, private training techniques create a spectrum of privacy—as determined by, for example, the level of noise introduced, or how many sensitive words are removed from the message—whereas E2EE is a binary classification.
@@
Private training techniques themselves provide a lens to illustrate the tension between E2EE and AI training: we can frame the guarantees of E2EE as “extreme” forms of various private training techniques, which correspond to degraded, impractical instantiations of each technique. For example, the guarantees of E2EE can be cast as differential privacy with $\epsilon=0$ (i.e., when the statistical distance between training data with or without a particular item included is 0). However, this value can only be achieved when private data is removed from training to begin with [wutschitz2023rethinking], at which point differential privacy is vacuous. Similarly, the guarantees of E2EE can be described as data sanitization in the case when all parts of an input are considered sensitive. Thus, sanitization techniques would remove E2EE messages entirely, which is equivalent to not training at all to begin with. As these examples show, to meet E2EE security, private training techniques essentially degrade to cases where training data is removed altogether.

### v2

@@ 4.1.1 AI Assistants Interacting with E2EE Data
As discussed in Section 2.2, a model’s training data consists of vast amounts of information which can be continuously supplemented by data collected from users (e.g., queries, usage patterns) after the model is deployed and used for inference. This data would necessarily be plaintext as a model cannot be trained on encrypted text. As such, if user data is collected and incorporated into the training process, the model is inherently, itself, a derivative of E2EE content.
Additionally, models have been observed to reproduce portions of the training data in their outputs, at times producing significant portions of training data, sometimes verbatim or nearly verbatim [cooper2024files, petroni2019languagemodelsknowledgebases, mccoy2023much]. The reasons for this phenomenon are not fully understood, but it is believed by many experts to be inherent in current machine learning and AI development techniques. Training on E2EE content thus raises the security threat of models exposing sensitive information from the training data even during regular (non-adversarial) use, constituting leaks of user data.
The memorization and reproduction of training data in models also introduces the risk of individuals exploiting this behavior to extract certain data points and patterns. Prior work has shown that adversarial queries can expose information in the training data, which would directly compromise the confidentiality of E2EE data [nasr2023scalable]. Individuals may potentially perform adversarial attacks, involving using the model (often in a black-box manner) as a means to acquire private data. Adversarial attacks are an active research area, and come in various forms, including membership inference attacks, where an adversary can determine whether a particular data point is part of the model’s training set [shokri2017membership, hu2022membership]; and training data extraction attacks, where an adversary can recover elements of the model’s training data often by generating candidate data points and using membership inference to verify whether the items are in the training set [274574].
@@ 4.1.2 Key Considerations
Given that integrated AI assistants may process potentially significant volumes of E2EE data, they are not an example of caveat features (as defined in Section 3.2). How integrated AI assistants are implemented in an E2EE application directly impacts how strongly an application preserves E2EE, and the addition of integrated AI may demote an application’s categorization within our taxonomy (Section 3.2). The risks outlined above underscore the following critical considerations for the implementation of AI assistants.
@@
    Which parties have access to the inputs and outputs from the model? Intermediaries that process plaintext user queries before relaying them to the model provider may have access to inputs and outputs from the model, in the form of chat logs. To preserve the guarantees of E2EE, inputs and outputs that depend on user messages should only be visible to the intended recipients of the messages.
@@ 4.2 How Privacy Technologies Can and Cannot Help
The privacy and confidentiality concerns outlined in the preceding subsection are by now well known to the machine learning, privacy, and security communities, and there exist various techniques to enhance privacy (according to various metrics and definitions) during inference and training.
Next, we evaluate whether the types of privacy enhancements that existing techniques offer meet E2EE confidentiality, considering a range of possible implementations. We find that existing privacy-enhancing techniques cannot address the security concerns we raise around integrating AI assistants with E2EE applications. We consider the following privacy-enhancing techniques: endpoint-local training and inference, fully homomorphic inference (FHE), trusted execution environments (TEEs), per-user models, data sanitization, differential privacy, federated learning, and multi-party computation (MPC). While some of the techniques and implementations that we consider reflect existing practices, others represent novel configurations.
Underlying our evaluation is a generalizable analytical framework designed to identify the key design considerations for implementing AI assistants in E2EE environments, structured around four key considerations which are detailed below. Our framework can be used to evaluate and compare different implementations of AI assistants with (or without) diverse privacy technologies, and the degree to which they uphold E2EE security, generalizing beyond the specific technologies that we examine in this paper. The four key framework considerations are interdependent: at each consideration point, design choices can either affirm the configuration’s compliance with E2EE or breach it (thereby rendering further considerations irrelevant). Figure 5 summarizes our evaluation framework as a flowchart, and Table 4.2 summarizes our analysis of various AI assistant implementations’ compatibility with E2EE under our framework. In Appendix A, we present an extended framework for evaluating the compatibility of arbitrary features (as opposed to just AI assistants) with E2EE, which generalizes the framework from this section and encompasses three levels: technical implementation, data inputs, and user control.
@@
The first design parameter is the location where data is processed, i.e., where the model is hosted. We distinguish between endpoint-local and non-endpoint-local designs, using the definitions from Section 3.1.
@@
Using endpoint-local models is not simple, however, as most consumer grade devices are unable to support the hefty computations involved in running an AI assistant. Local processing is feasible in limited circumstances: e.g., in certain use cases of Apple Intelligence and Samsung’s Galaxy AI (see Section 2.5). But even with custom hardware, not all AI assistant queries can be supported with the limited computational faculties of a smartphone, and many features require sending plaintext messages off the device.
When non-endpoint-local models are used to answer queries, the compatibility with E2EE depends on Considerations #2 and #3 below.
Consideration #2: If the model is non-endpoint-local, is its processing of user queries privacy-enhanced, or in plaintext?
For non-endpoint-local models, the confidentiality of data depends on whether processing of queries occurs in plaintext or in some privacy-enhanced form (i.e., encrypted or contained in secure hardware) on the servers. Plaintext inference and training is not compatible with E2EE, as they would expose user data to third parties. Privacy-enhanced processing provides additional protection. There are various technologies that could implement some kind of “privacy enhancement”—such as fully homomorphic encryption (FHE), and trusted execution environments (TEEs)—which differ on their confidentiality guarantees, performance, and other practical considerations. The compatibility of privacy-enhanced processing with E2EE depends on the nature of the confidentiality offered by the technology being used, which is the focus of Consideration #3.
Consideration #3: If the model is non-endpoint-local and performs privacy-enhanced processing, what is the type of the confidentiality guarantees?
As discussed in Section 2.4.2, different security technologies—such as TEEs and encryption—provide different types of confidentiality under different conditions. Established cryptographic techniques tend to provide the same type of security as E2EE under types of conditions, and are therefore more likely to be compatible with E2EE.
Fully homomorphic encryption (FHE) provides an instructive example. FHE enables arbitrary computations to be performed on encrypted data without the need to decrypt the data, and yields the result of the computation still in encrypted form. FHE makes it possible to “outsource” a computation on private data without revealing anything about the data to the outsourcing provider. FHE would thus allow non-endpoint-local models to process user messages for inference without decrypting them—at least in theory.
@@
As described in Section 2.4.2, hardware-backed security is different from E2EE security. Hardware solutions represent a substantial security/privacy improvement over plaintext inference, which may well be a fruitful avenue for exploration, beyond the scope of this paper. Our scope is to stress that the importance of clearly distinguishing them from E2EE and cryptographic security guarantees.
@@
Any model training introduces an additional critical consideration: who can query models trained with whose E2EE data? Since a model is a derivative of its training data, it is important that queries over this training data are consistent with the guarantees of E2EE.
@@
There are two possible configurations for this design choice: shared or per-user models, as defined in Section 3.1. Shared models are not compatible with E2EE, since users would be able to receive inference outputs that are a function of other users’ E2EE data, which violates E2EE confidentiality. In practice, this means that E2EE data may be exposed to entities that can query the model (e.g., due to the possibility of leakage of training data and adversarial attacks). As we discuss in more detail in Section 4.2.1, mitigations for adversarial attacks (e.g., differential privacy) are insufficient to fully address this concern. Shared models are incompatible with training on E2EE content, irrespective of the implementation.
@@
On the other hand, per-user models can be compatible with E2EE, provided that applications enforce strict isolation between user data, and that any components that depend on E2EE data must themselves be adequately privacy-enhanced in the sense outlined in Considerations #2 and #3. There are a number of ways an application could support per-user models. For example, a service could maintain a separate copy of a foundation model for each user in the system, which is fine-tuned with that user’s data. This approach would maintain strict isolation, but require large amounts of resources for applications with many users. A more practical approach is to use retrieval augmentation techniques: the application hosts a single foundation model, which is not trained on user data, but is dynamically refined during inference using user-dependent data sources [martineau_what_2023]. [^37] See [wutschitz2023rethinking] for more on retrieval augmentation. The application could store E2EE data for every user locally (e.g., [martineau_what_2023]), and retrieve and incorporate it locally during inference—temporarily fine-tuning the model. After the query is processed, the base model should revert to its prior state, which was not dependent on any E2EE data, resulting in a stateless augmentation process. While this reversion or statelessness would largely address the specific concern of Consideration #4—namely, which users can query the models trained with the E2EE data of which users—the storage and processing of the per-user E2EE data and the temporarily fine-tuned model (which is itself E2EE data) would, in turn, need to be scrutinized under Considerations #1–#3 and comply with the recommendations we put forward throughout this paper.
Applying this framework
A specific implementation of an AI assistant can be evaluated with respect to each of the four design choices in our framework. From the results of this evaluation, we can determine the implementation’s compatibility with E2EE. For example, the analysis of local inference would be as follows:
(1) local, (2) N/A, (3) N/A, (4) N/A, therefore fully compatible with E2EE.
Example analyses of a range of possible implementations are presented in Table 4.2.
@@
l—llll Implementation Location Privacy-enhanced Guarantees Shared or per-user
@@
Table 1: Example analyses of different implementations of AI assistants under our framework. Green tags denote an implementation compatible with E2EE, yellow tags denote an implementation that is not compatible with E2EE but has additional protections, and red tags denote implementations that are most vulnerable.
Connection to our taxonomy
Recall that Section 3.2 describes a taxonomy from “Strict E2EE” to “No E2EE.” The (in)compatibility of an AI assistant implementation with E2EE, as assessed under the evaluation framework just presented, directly impacts an application’s categorization under our taxonomy. An application’s original features, before integrating an AI assistant, determine what we call its initial categorization; if AI features are added, a fresh evaluation and categorization are necessary, and the new categorization will depend on the implementation details highlighted in our evaluation framework.
The addition of a non-E2EE-compatible AI assistant implementation to an application that is initially in Category 1 or 2 will result in the application being “demoted” out of those two categories, as the added AI feature will render E2EE content no longer fully protected in the way that E2EE promises. In more detail:
  • If the AI assistant implementation is fully compatible with E2EE (green tags in Table 4.2), the application maintains its initial category (from before the AI feature was added).
  • If the AI assistant implementation is not fully compatible with E2EE (yellow or red tags in Table 4.2), the application is demoted to Category 3, 4, or 5, as follows: If the AI assistant is off by default or provides an option to turn the AI assistant off application-wide, the application is demoted to Category 3. If the AI assistant is on by default and only provides an option to turn off in specific contexts (e.g., for individual chats), the application is demoted to Category 4. If the AI assistant cannot be turned off, the application is demoted to Category 5.
4.2.1 Discussion on Privacy-Preserving Training
This section considers a range of privacy-preserving training methods—differential privacy, data sanitization, federated learning, and multi-party computation—and discusses how the types of privacy guarantees they provide are different from E2EE confidentiality. None of these techniques can guarantee E2EE confidentiality for training shared models on E2EE content.
@@ 4.2.1 Discussion on Privacy-Preserving Training
Differential privacy (DP) is a method for designing algorithms to compute statistics over datasets such that the results do not depend “too much” on any individual data entry. In contrast with encryption’s binary security guarantee, DP is parameterized by an “acceptable” amount of “privacy loss,” representing dependence upon individual data entries (encoded via a parameter $\varepsilon$)  [DBLP:conf/tamc/Dwork08]. Applying DP to machine learning is an active area of research; training with DP incurs a computational overhead and loss in accuracy compared to a non-DP model [nist_dp_ml, Ponomareva_2023]. Though DP is often discussed alongside cryptographic techniques, DP offers a categorically different security guarantee than encryption, applicable under different conditions.
@@
Data sanitization aims to remove “sensitive” information from data that is aggregated into training datasets. This presents the difficulty of determining what information is “sensitive;” sensitivity depends on context and there is no generally correct heuristic for this procedure [brown2022does]. In contrast, the strict confidentiality of encryption is agnostic to message contents, and provides systemic security benefits by protecting mundane communications alongside sensitive ones (as discussed in Section 2.1.3).
Federated learning is another approach intended to enhance the privacy of training data, by having end-users locally train local models with their personal data. The resulting local model information is then sent to and aggregated into a single model by a central entity. The privacy benefit is that the raw personal data does not leave users’ devices. But the local models, which are dependent on that personal data, do leave users’ devices. In turn, the final centralized model is still dependent on users’ personal data as it depends on these local models. Thus, federated learning does not address the issues of training data leakage through model use or adversarial attacks.
Multi-party computation (MPC) is a technique that allows multiple parties (e.g., users or companies), each with a private data set, to collectively compute a function (e.g., model training) on all of their data sets combined and access the result (e.g., a trained model), without directly sharing their data sets with each other. To use MPC, all parties must agree on a function they want to compute. MPC provides the guarantee that parties will learn the desired function output (e.g., the trained model) and nothing more than that output. [^38] Necessarily, by learning the output, parties may also learn information that can be inferred from that output. MPC does not constrain the function parties choose to compute, and in particular, allows for the function output to be dependent on, and reveal information about, the input data sets—such as when the output is a trained model. [^39] The confidentiality guarantee offered by MPC thus depends heavily on the function that the parties choose to compute, which they can choose freely. For example, if the function they choose to compute is the identity function—i.e., the output is simply a copy of all of the parties’ data sets—then MPC provides no confidentiality at all. If the chosen function consists of aggregate statistics—e.g., the average salary of the many employees in an industry—then MPC provides a stronger guarantee. A trained model resulting from an MPC computation is just as dependent on its training data as a model trained without MPC. Thus, MPC does not address the issues of training data leakage through model use or adversarial attacks.
All of the above techniques provide valuable privacy protections. However, their privacy goals are different in kind from the privacy goals of E2EE systems. As such, they cannot achieve E2EE confidentiality for training shared AI models on E2EE content. Future advances in the technologies discussed will likewise be valuable but will not bring these technologies closer to E2EE, because the type of privacy protection they provide is different in kind from E2EE confidentiality. Put differently, private training techniques create a spectrum of privacy—as determined by, for example, the level of noise introduced, or how many sensitive words are removed from the message—whereas E2EE is a binary kind of guarantee.
@@
Private training techniques themselves provide a lens to illustrate the tension between E2EE and AI training: we can frame the guarantees of E2EE as “extreme” forms of various private training techniques, which correspond to degraded, impractical instantiations of each technique. For example, the guarantees of E2EE can be cast as differential privacy with $\varepsilon=0$ (i.e., when the statistical distance between training data with or without a particular item included is 0). However, this value can only be achieved when private data is removed from training to begin with [wutschitz2023rethinking], at which point differential privacy is vacuous. Similarly, the guarantees of E2EE can be described as data sanitization in the case when all parts of an input are considered sensitive. Thus, sanitization techniques would remove E2EE messages entirely, which is equivalent to not training at all to begin with. As these examples show, to meet E2EE security, private training techniques essentially degrade to cases where training data is removed altogether.
4.3 How AI Features are Integrated Matters for E2EE
At time of writing, messaging applications with integrated AI assistants typically offer multiple ways of interfacing with AI features. These differences may have implications for E2EE compatibility. For example, at time of writing, WhatsApp offers at least three distinct ways to invoke Meta AI.
  1. 1.
    Through the search bar, which is labeled “Ask Meta AI or Search.” When the search bar is selected, it previews some example prompts for Meta AI. [^40] The search bar can be used either to locally search a user’s WhatsApp conversations or to “Ask Meta AI,” and appears at the top of a user’s conversation list.
  2. 2.
    Through a button with the Meta AI logo at the bottom right of the Chats page, which takes you to a one-on-one chat with Meta AI.
  3. 3.
    Through typing “@Meta AI” in a chat with other users. When typing the “@” symbol, “Meta AI” also appears as a recipient like any user in a group chat, and is put at the top of the list.
It is the third means of invoking Meta AI, which is positioned within the messaging service—and at times, seems to present Meta AI much like another “user” or “contact” within a user’s conversations and group chats—which falls squarely within the scope of our concerns and recommendations.
The first way of invoking Meta AI is a standalone AI feature (as defined in Section 2.2.4), because only user-initiated search-bar queries that explicitly invoke Meta AI are processed by Meta AI, without involving any E2EE content—and thus is outside the scope of our concerns.
The second way of invoking Meta AI is also, technically speaking, a standalone feature not integrated with E2EE content. This is because the only content accessible to the AI is in dedicated user-initiated interactions or “chats” between the user and Meta AI (and nobody else)—in effect, a chat-like user interface for a direct user-AI interaction that involves no AI processing of E2EE content, [^41] We consider content in E2EE conversations between users to be E2EE content, and distinguish this from direct interactions between a single initiating user and an AI assistant, which we do not consider to contain any E2EE content even if packaged in a chat-like interface. This is consistent with our definition of E2EE content in Section 3.1. just like the search bar. However, this chat-like interface may create potential for confusion. As noted in Section 2.2.4, standalone AI features are generally outside our scope, unless the provider creates the impression that the AI feature is itself E2EE. Because of ambiguities in the claims that Meta makes about whether “chats” with Meta AI are E2EE, the one-on-one chats with Meta AI are within the scope of our recommendations.424242 Section 7.2 provides more detailed discussion of the ambiguities.